Lorge Corporation has collected the following information after its first year of sales. Sales were $2,500,000 on 100,000 units; selling expenses $250,000 (40% variable and 60% fixed); direct materials $1,370,600; direct labor $250,000; administrative expenses $270,000 (20% variable and 80% fixed); and manufacturing overhead $322,000 (70% variable and 30% fixed). Top management has asked you to do a CVP analysis so that it can make plans for the coming year. It has projected that unit sales will increase by 10% next year.

Compute (1) the contribution margin for the current year and the projected year, and (2) the fixed costs for the current year. (Assume that fixed costs will remain the same in the projected year.)

(1) Contribution margin for current year

$

Contribution margin for projected year

$

(2) Fixed costs for current year

$

eTextbook and Media

  • eTextbook

Compute the break-even point in units and sales dollars for the first year. (Round contribution margin ratio to 2 decimal places e.g. 0.15 and final answers to 0 decimal places, e.g. 2,510.)

Break-even point units
Break-even point

$

eTextbook and Media

The company has a target net income of $150,000. What is the required sales in dollars for the company to meet its target?

Sales dollars required for target net income

$

eTextbook and Media

If the company meets its target net income number, by what percentage could its sales fall before it is operating at a loss? That is, what is its margin of safety ratio? (Round answer to 1 decimal place, e.g. 10.5.)

Margin of safety ratio %

eTextbook and Media

The company is considering a purchase of equipment that would reduce its direct labor costs by $110,000 and would change its manufacturing overhead costs to 30% variable and 70% fixed (assume total manufacturing overhead cost is $322,000, as above). It is also considering switching to a pure commission basis for its sales staff. This would change selling expenses to 90% variable and 10% fixed (assume total selling expense is $250,000, as above). Compute (1) the contribution margin and (2) the contribution margin ratio, and recompute (3) the break-even point in sales dollars. (Round contribution margin ratio to 2 decimal places, e.g. 25.25 and all other answers to 0 decimal places, e.g. 2,520. Use the current year numbers for calculations.)

1. Contribution margin

$

2. Contribution margin ratio %
3. Break-even point

$
Answer #1

1)

Current year Variable Fixed
Direct material 1370600
Direct labor 250000
overhead 322000*70%=225400 322000*30%= 96600
selling 250000*40%=100000 250000*60%=150000
Administrative cost 270000*20%=54000 270000*80%= 216000
Total 2000000 462600
selling price per unit 2500000/100000= 25 per unit
Variable cost per unit 2000000/100000=$20per unit
contribution per unit 5 per unit

a)contribution margin for current year =sales -variable cost

                       = 2,500,000 - 2,000,000

                      = 500,000

b)contribution margin for projected year:units sales *contribution per unit

                                                          = 110000*5

                                                           = $ 550000

#increased unit sales :110000 units       [100000(1+.10)]

c)Fixed cost for current year = $ 462600

d)Break even point in units =Fixed cost /contribution per unit

                         = 462600/5

                          = 92520 units

Breakeven point in $ = Fixed cost /contribution margin ratio

                   = 462600/.20

                  = $ 2313000

**Contribution margin ratio =contribution /price

                            = 5/25

                             = .20 or 20%

e) Sales dollars required for target net income =[Fixed cost+Target income ]/contribution margin ratio

                        = [462600+150000]/.20

                     = 612600/.20

                  = $ 3063000

f)Margin of safety ratio = [Target sales-Breakeven point sales ]/Target sales

              =[3063000-2313000]/3063000

              = 750000/3063000

              = 24.5%

2).

Current year Variable Fixed
Direct material 1370600
Direct labor 250000 -110000 = 140000
overhead 322000*30%=96600 322000*70%= 225400
selling 250000*90%=225000 250000*10%=25000
Administrative cost 270000*20%=54000 270000*80%= 216000
Total 1886200 466400
selling price per unit 2500000/100000= 25 per unit
Variable cost per unit 1886200/100000=$18.862per unit
contribution per unit 5 per unit

a)contribution margin =sales -variable cost

                       = 2,500,000 - 1886200

                      =613800

b)contribution margin ratio = 613800/2500000

                   = 24.55%

c)Breakeven point ($) =Fixed cost /contribution margin ratio

          = 466400 /.2455

             = $ 1,899,796